The next number in the series
41 37 38 19 15 16 8 is

Answer by Theo(13342)

looks like the pattern is - 4 then + 1 then / 2.

41 - 4 = 37
37 + 1 = 38
38 / 2 = 19

19 - 4 = 15
15 + 1 = 16
16 / 2 = 8

if the pattern holds, the then the next 3 numbers will be:

8 - 4 = 4 *****
4 + 1 = 5
5 / 2 = 2.5

***** the next number in the seqeuence will be 4 if the pattern holds.